Please help me find x in 13x^2-650x-6000=0. I have tried many combinations using trial and error method.
Answer by mukhopadhyay(490)

13x^2-650x-6000=0
=> 13(x^2-50x) = 6000
=> 13(x^2-50x+25^2) = 6000 + 13*25^2
=> 13(x-25)^2 = 6000+8125
=> 13(x-25)^2 = 14125
=> (x-25)^2 = 14125/13 = 25*565/13
=> (x-25) = 5*sqrt(565/13) or x = -5*sqrt(565/13)
=> x = 25 + 5*sqrt(565/13) or x = 25 - 5*sqrt(565/13)